# Random Pareto front surfaces

### v1

@@ 5.4 Air pollution example
The previous examples shown in this work have largely been focussed on static problems which have no time dependency. Nevertheless, all of the ideas that we have presented so far can also accommodate for dynamic problems, where there is indeed some time dependency. As an illustrative example, we will now consider how these ideas can be used on a real-world time series data set. Precisely, we study the air pollution data obtained at the North Kensington (UKA00253) air monitoring station [^10] We have curated this data from the UK-AIR website (https://uk-air.defra.gov.uk/), which is a publicly accessible domain that is ran by the Department for Environment, Food & Rural Affairs (DEFRA) in the United Kingdom. in west London.
@@ 6 Conclusion
In this work we have studied the space of Pareto front surfaces. We presented a novel result which showed that any Pareto front surface can be explicitly parametrised using polar coordinates. We then used this polar parametrisation result in order to establish many general concepts over the space of Pareto front surfaces. Notably, we focussed our attention on the stochastic setting where the Pareto front surface itself was random. We showed that this random surface admits a stochastic process representation (10), based on our polar parametrisation, which makes it amenable to standard ideas from functional data analysis. In particular, we used this stochastic representation in order to define and compute statistics for random Pareto front surfaces. We then demonstrated the usefulness of these new probabilistic ideas on some interesting applications such as: visualisation, uncertainty quantification and extreme value theory. In the following paragraphs, we highlight some interesting directions for future research.
@@
Throughout this work, we have focussed our attention on the interpolated Pareto front (Definition 2.4), which was defined using the weak Pareto partial ordering. This type of interpolation is sensible because it is consistent with the weak Pareto partial ordering on sets (Definition 3.2): $\mathcal{Y}_{\boldsymbol{\eta}}^{\text{int}}[A]\preceq\mathcal{Y}_{\boldsymbol{\eta}}^{\text{int}}[B]$ if $A\subseteq B\subseteq\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$. In practice, one might also consider using alternate interpolation (or approximation) schemes such as the ones based on Delaunay triangulation (Hartikainen et al., 2012), sandwiching (Bokrantz and Forsgren, 2013) or hyperboxing (Dächert and Teichert, 2020; Eichfelder and Warnow, 2022). Notably these other schemes might be beneficial when we are interested in approximating a Pareto front of a continuous set $Y\subset\mathbb{R}^{M}$ using a finite subset $\hat{Y}\subseteq Y$. In this setting, the interpolated weak Pareto front of $\hat{Y}$ would always give us a lower bound estimate of the actual front. In contrast, a different interpolation scheme might be able to give us a more accurate approximation of the target front, with respect to some frontier loss (Section 3.3). Naturally, many of the results which we have shown here continues to hold even when we change the interpolation scheme. The only notable difference is that the corresponding projected length function $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}$, which arises in the polar parametrisation (Theorem 3.1), might no longer be written in terms of the length scalarisation function (2). Instead, the new projected length function would likely depend on a different scalarisation function—one which accurately caters for the interpolation scheme at hand. We leave the analysis of these other interpolation schemes for future research.
Constrained regression
@@
Our polar parametrisation result implies that any Pareto front surface is completely characterised by its projected length function. Whilst our result in Proposition 3.1 gives two necessary and sufficient conditions that a length function must satisfy in order for it to induce a valid Pareto front surface. An interesting direction for future work would be focussed on learning this projected length function directly from observational data. Note that the first constraint (C1), positivity, can be easily handled by just considering the logarithm of the length function instead. Whilst the second constraint (C2), the maximum ratio condition, is much less trivial to work with. Specifically, by taking the logarithm and rearranging the terms, we see that this latter condition is very similar to a Lipschitz type constraint, which is known to be challenging to work with.
Advancing the theory
Many of the ideas that we have established in this work can be generalised and adapted even further. For instance, we only exploited a few basic ideas from functional data analysis in order to define the Pareto front surface statistics in Section 4. Other ideas such as functional regression and functional principal component analysis were not explored in any detail and could be investigated in the future. Similarly, in Section 5.3, we gave a proof of concept on how one could extend well-known results from extreme value theory to work in the multivariate setting under the Pareto maximum. A practical direction for future work could focus on pushing these preliminary ideas even further.
@@ Appendix A Proofs
Equipped with Assumptions 4.1 and 4.2, we will show that the expectation in (12) satisfies the two conditions in Proposition 3.1 and therefore is a valid Pareto front surface. Firstly, C1 is satisfied because the lengths are assumed to be positive almost surely, which implies that $\mathbb{E}_{\boldsymbol{\omega}}[\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega})]]>0$ for all $\boldsymbol{\lambda}\in\mathbb{R}^{M}$. Secondly, C2 is satisfied because $Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega})\in\mathbb{L}_{\boldsymbol{\eta}}$ is a Pareto front surface almost surely and therefore it satisfies C2 almost surely,
@@
A.9 Proof of Theorem 5.1
See for instance Leadbetter et al. (1983, Theorem 1.4.2).
$\blacksquare$
A.10 Proof of Theorem 5.2
See for instance Leadbetter et al. (1983, Theorem 1.6.2).
$\blacksquare$
A.11 Proof of Proposition 5.2

### v2

@@ 5.4 Air pollution example
All of the ideas that we have presented so far can also accommodate for dynamic problems, where the Pareto front surface distribution can vary in time. As an illustrative example, we will now consider how these ideas can be used on a real-world time series data set. Precisely, we study the air pollution data obtained at the North Kensington (UKA00253) air monitoring station [^9] We have curated this data from the UK-AIR website (https://uk-air.defra.gov.uk/), which is a publicly accessible domain that is ran by the Department for Environment, Food & Rural Affairs (DEFRA) in the United Kingdom. in west London.
@@ 6 Conclusion
In this work we presented a novel result which showed that any Pareto front surface can be explicitly parameterised using polar coordinates. We exploited this polar parameterisation result in order to define and compute statistics for random Pareto front surfaces. We then demonstrated the usefulness of these new probabilistic ideas on some interesting applications such as: visualisation, uncertainty quantification and extreme value theory. In the following paragraphs, we highlight some interesting directions for future research.
@@
Throughout this work, we have focussed our attention on the interpolated Pareto front (Definition 2.4), which was defined using the weak Pareto partial ordering. This type of interpolation is sensible because it is consistent with the weak Pareto partial ordering on sets (Definition 3.2): $\mathcal{Y}_{\boldsymbol{\eta}}^{\text{int}}[A]\preceq\mathcal{Y}_{\boldsymbol{\eta}}^{\text{int}}[B]$ if $A\subseteq B\subseteq\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$. In practice, one might also consider using alternate interpolation (or approximation) schemes such as the ones based on Delaunay triangulation (Hartikainen et al., 2012), sandwiching (Bokrantz and Forsgren, 2013) or hyperboxing (Dächert and Teichert, 2020; Eichfelder and Warnow, 2022). Notably these other schemes might be beneficial when we are interested in approximating a Pareto front of a continuous set $Y\subset\mathbb{R}^{M}$ using a finite subset $\hat{Y}\subseteq Y$. In this setting, the interpolated weak Pareto front of $\hat{Y}$ would always give us a lower bound estimate of the actual front. In contrast, a different interpolation scheme might be able to give us a more accurate approximation of the target front, with respect to some frontier loss (Section 3.3). Naturally, many of the results which we have shown here continues to hold even when we change the interpolation scheme. The only notable difference is that the corresponding projected length function $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}$, which arises in the polar parameterisation (Theorem 3.1), might no longer be written in terms of the length scalarisation function (2). Instead, the new projected length function would likely depend on a different scalarisation function—one which accurately caters for the interpolation scheme at hand. We leave the analysis of these other interpolation schemes for future research.
Modelling the Pareto front surface
@@
Our polar parameterisation result implies that any Pareto front surface is completely characterised by its projected length function. Proposition 3.1 gives two necessary and sufficient conditions that a length function must satisfy in order for it to induce a valid Pareto front surface. An interesting direction for future work would be focussed on learning this projected length function directly from observational data. Note that the first constraint (C1), positivity, can be easily handled by just considering the logarithm of the length function instead. Whilst the second constraint (C2), the maximum ratio condition, is much less trivial to worth with. Specifically, by taking the logarithm and rearranging the terms, we see that this latter condition is very similar to a Lipschitz type constraint, which can be a difficult constraint to enforce in practice (Virmaux and Scaman, 2018).
@@ Appendix A Proofs
Equipped with footnotes 5 and 4.2, we will show that the expectation in (12) satisfies the two conditions in Proposition 3.1 and therefore is a valid Pareto front surface. Firstly, C1 is satisfied because the lengths are assumed to be positive almost surely, which implies that $\mathbb{E}_{\boldsymbol{\omega}}[\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega})]]>0$ for all $\boldsymbol{\lambda}\in\mathbb{R}^{M}$. Secondly, C2 is satisfied because $Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega})\in\mathbb{L}_{\boldsymbol{\eta}}$ is a Pareto front surface almost surely and therefore it satisfies C2 almost surely,
@@
A.9 Extreme value theory
We begin this subsection by first recalling the statement of Fisher–Tippett–Gnedenko theorem (Fisher and Tippett, 1928; Gnedenko, 1943) and the Balkema–de Haan–Pickands theorem (Balkema and de Haan, 1974; Pickands, 1975) in Theorem A.1 and Theorem A.2, respectively. The proof of these two results can be found in standard textbooks; for example, see the proofs presented by Leadbetter et al. (1983, Theorem 1.4.2) and Leadbetter et al. (1983, Theorem 1.6.2), respectively. Afterwards, we move on to Section A.9.1, where we present the proof of Proposition 5.2.
Definition A.1 (Maximum domain of attraction)
Theorem A.1 (Extreme value distribution)
Definition A.2 (Conditional excess distribution)
$F_{u}(x):=\mathbb{P}[Y-u\leq x|Y>u]=\frac{F(u+x)-F(u)}{1-F(u)},$
Theorem A.2 (Generalised Pareto distribution)
A.9.1 Proof of Proposition 5.2